OpenStudy (giitter):

Simplify: (2+1)^2 @Unavailabilityy

OpenStudy (giitter):

Will medal and fan if helpful ^_^

jhonyy9 (jhonyy9):

2+1=3 3^2 = ?

OpenStudy (unavailabilityy):

Okay so This is pretty easy Just follow PEMDAS Parenthesis goes first 2+1 is three Three to the second power means 3 times 3. What is your answer?

jhonyy9 (jhonyy9):

3^2 = 3*3 = ?

OpenStudy (giitter):

6?

OpenStudy (unavailabilityy):

No. You added. Exponents is repeated multiplication.

OpenStudy (unavailabilityy):

What is three times three?

OpenStudy (giitter):

So 9? would you want the answers that are there 3 5 6 9

OpenStudy (unavailabilityy):

You are correct The answer is 9. :)

HanAkoSolo (jamierox4ev3r):

@Unavailabilityy Did a really really good job! The key to remember with problems like these is that you want to solve whatever is in parentheses first before you perform any other operations on it

HanAkoSolo (jamierox4ev3r):

@GIitter

OpenStudy (unavailabilityy):

:)

HanAkoSolo (jamierox4ev3r):

And just in case you'd like some clarification on what PEMDAS is, here you go: 

OpenStudy (giitter):

THANK YOU!